# Universal Hitchin moduli spaces

### v1

@@ 2.3. Harmonicity equations and hyperkähler structure
where $A$ is a connection on $E_{K}$ and $\psi\in\Omega^{1}(X,E_{K}(\mathfrak{k}))$. This simple fact induces a one-to-one correspondence
(2.7) ${\curly A}\times\Omega^{1}(X,E_{K}(\mathfrak{k}))$ $\longrightarrow$ ${\curly D}$
$(A,\psi)$ $\longmapsto$ $D:=A+i\psi$
@@
(2.8) $\mathbf{I}(a,\dot{\psi})$ $=(Ja,-J\dot{\psi})$
(2.9) $\mathbf{J}(a,\dot{\psi})$ $=(-\dot{\psi},a)$
(2.10) $\mathbf{K}(a,\dot{\psi})$ $=(-J\dot{\psi},-Ja),$
@@
(2.14) $\begin{array}[]{l}F_{A}-\frac{1}{2}[\psi,\psi]=0,\\ d_{A}\psi=0,\\ d_{A}^{\ast}\psi=0.\end{array}$
@@
The harmonicity condition for the reduction $h$ admits an interpretation in terms of the the holonomy representation $\rho:\pi_{1}(X)\to G$ corresponding to $D$. The reduction $h$ is equivalent to a $\pi_{1}(X)$-equivariant smooth map
$\tilde{h}:\tilde{X}\longrightarrow G/K,$
where $\tilde{X}$ is the universal covering of $X$. Here $\pi_{1}(X)$ acts on $\tilde{X}$ by Deck transformations, and via the representation $\rho$ on $G/K$. The symmetric space $G/K$ is equipped with a canonical $G$-invariant Riemannian metric determined by the Killing form $B$, and (2.14) is equivalent to the harmonicity of $\tilde{h}$ in the usual sense.
@@
The following theorem, due to Donaldson [Do] for $G=\mathrm{SL}(2,\mathbb{C})$, and Corlette [C] in general, is one of the main structural results of the theory.
@@
2.4. $\boldsymbol{G}$-Higgs bundles and Hitchin equations
@@
There are appropriate notions of stability, semistability and polystability and we can consider $\mathcal{M}(G)$ to be the moduli space of isomorphism classes polystable $G$-Higgs bundles over $X$.
Let $h\in\Omega^{0}E(X,G/K)$ be a smooth reduction of structure group of $E$ to $K$ and let $E_{h}$ be the corresponding smooth principal $K$-bundle. Let $F_{h}\in\Omega^{2}(X,E_{h}(\mathfrak{k}))$ be the curvature of the unique connection compatible with $h$ and the holomorphic structure on $E$ (defined by the Chern–Singer correspondence). Here $E_{h}(\mathfrak{k})$ is the adjoint bundle of $E_{h}$.
@@
$\tau\colon\Omega^{1,0}(X,E(\mathfrak{g}))\to\Omega^{0,1}(X,E(\mathfrak{g}))$
@@
A reduction $h$ of structure group of $E$ to $K$ satisfies the Hitchin equation
@@
(2.15) $\begin{array}[]{l}F_{A}-[\varphi,\tau(\varphi)]=0\\ \bar{\partial}_{A}\varphi=0.\end{array}$
Here $\bar{\partial}_{A}$ is the $(0,1)$ part of the covariant derivative $d_{A}$ defined by $A$. This is of course the Dolbeault operator defined by the holomorphic structure $J_{A}$ on $E_{G}$ corresponding to $A$ via de Chern–Singer correspondence. Here $E_{G}$ is the smooth principal $G$-bundle obtained frm $E_{K}$ by extension of structure group. The gauge group $\curly K$} of automorphisms of $E_{K}$ acts on the space of solutions and the moduli space of solutions is
@@
$\mathcal{M}(G)\cong\mathcal{M}^{\operatorname{Hit}}(G)$
To explain this correspondence we interpret the moduli space of $G$-Higgs bundles in terms of pairs $(J_{E},\varphi)$ consisting of a holomorphic structure $J_{E}$ on the smooth $G$-bundle $E_{G}$ obtained from $E_{K}$ by the extension of structure group, and $\varphi\in\Omega^{1,0}(X,E_{G}(\mathfrak{g}))$ satisfying $\bar{\partial}_{E}\varphi=0$. Such pairs are in correspondence with $G$-Higgs bundles $(E,\varphi)$, where $E$ is the holomorphic $G$-bundle defined by $J_{E}$ on $E_{G}$, and $\bar{\partial}_{E}\varphi=0$, that is $\varphi\in H^{0}(X,E(\mathfrak{g})\otimes K_{X})$. The moduli space of polystable $G$-Higgs bundles $\mathcal{M}(G)$ can now be identified with the orbit space
@@
2.5. Hitchin equations and hyperkähler structure
@@
The Chern–Singer correspondence [Sin, W] establishes an isomorphism
(2.16) ${\curly A}$ $\longrightarrow$ ${\curly C}$
(2.17) $A$ $\longmapsto$ $J_{A}$
@@
(2.18) $\Omega^{0,1}(X,E_{G}(\mathfrak{g}))$ $\longrightarrow$ $\Omega^{1}(X,E_{K}(\mathfrak{k}))$
(2.19) $\alpha$ $\longmapsto$ $a:=\alpha-\tau(\alpha)$
@@
(2.20) ${\curly A}$ $\longrightarrow$ $\Omega^{2}(X,E_{K}(\mathfrak{k}))$
(2.21) $A$ $\longmapsto$ $F_{A}.$
@@
(2.22) $\begin{array}[]{ccc}{\curly A}\times\Omega&\longrightarrow&\Omega^{2}(X,E_{K}(\mathfrak{k}))\\ (A,\varphi)&\mapsto&F_{A}-[\varphi,\tau(\varphi)].\end{array}$
@@
$\begin{array}[]{ccc}J_{1}(\alpha,\eta)&=&(i\alpha,i\eta)\\ J_{2}(\alpha,\eta)&=&(-i\tau(\eta),i\tau(\alpha))\\ J_{3}(\alpha,\eta)&=&(\tau(\eta),-\tau(\alpha)),\end{array}$
These complex structures correspond to the complex structures $\mathbf{I},\mathbf{J}$ and $\mathbf{K}$ in (2.8) via the the usual Chern–Singer correspondence.
@@
We thus have that ${{\boldsymbol{\mu}}}^{-1}({0})/{\curly K}$ is the moduli space $\mathcal{M}^{\operatorname{Hit}}(G)$ of solutions to Hitchin equations (2.15). In particular, if we consider the set ${{\boldsymbol{\mu}}}_{*}^{-1}({0})$ of irreducible solutions (equivalently, smooth) one has that
@@
is a hyperkähler manifold which, by Theorem 2.7, is homeomorphic to the subvariety of smooth points of the moduli space $\mathcal{M}(G)$, consisting of stable and simple $G$-Higgs bundles on $E_{G}$.
@@ 2.6. Non-abelian Hodge correspondence
$\begin{array}[]{ccc}{\curly A}\times\Omega&\longrightarrow&{\curly D}\\ (A,\varphi)&\longmapsto&A-i\varphi+i\tau(\varphi)\end{array}$
@@
Now, Theorems 2.7 and 2.5 can be regarded as existence theorems, establishing the non-emptiness of the hyperkähler quotient, obtained by focusing on different complex structures. For Theorem 2.7 one gives a special status to the complex structure $J_{1}$. Combining the symplectic forms determined by $J_{2}$ and $J_{3}$ one has the $J_{1}$-holomorphic symplectic form $\omega_{c}=\omega_{2}+i\omega_{3}$ on ${\curly A}\times\Omega$. The gauge group $\curly{G}={\curly K}^{\mathbb{C}}$ acts on ${\curly A}\times\Omega$ preserving $\omega_{c}$. The symplectic quotient construction can also be extended to the holomorphic situation (see e.g. [K]) to obtain the holomorphic symplectic quotient $\{(J_{E},\varphi)\;:\;\bar{\partial}_{E}\varphi=0\}/\curly{G}$. What Theorem 2.7 says is that for a class $(J_{E},\varphi)]$ in this quotient to have a representative (unique up to $K$-gauge) satisfying $\mu_{1}=0$ it is necessary and sufficient that the pair $(J_{E},\varphi)$ be polystable. This identifies the hyperkähler quotient to the set of equivalence classes of polystable $G$-Higgs bundles on $E_{G}$. If one now takes $J_{2}$ on ${\curly A}\times\Omega$ or equivalently $\curly D$} with $\mathbf{J}$ and argues in a similar way, one gets Theorem 2.5 identifying the hyperkähler quotient to the set of equivalence classes of reductive flat connections on $E_{G}$.
@@ 3.1. Kähler fibrations
In this section we recall some basic aspects of the theory of Kähler fibrations, which we will use later. We follow closely [GLS, M].
A Kähler fibration is a holomorphic fibre bundle $\curly{X}\to\curly{J}$, with typical fiber a Kähler manifold $({\curly D},\omega_{\curly D})$, and such that transition functions between local holomorphic trivializations are contained in the group of Kähler isometries of $\curly D$}. Equivalently, $\curly{X}$ admits a smoothly varying Kähler structure on the fibres, that we shall denote $\widehat{\boldsymbol{\omega}}$. An Ehresmann connection $\Gamma$ on $\curly{X}$, given by a distribution of horizontal subspaces
@@
The following basic result shows that the existence of closed $(1,1)$-form restricting on the fibres to $\widehat{\boldsymbol{\omega}}$ is indeed a sufficient condition for the existence of a Kähler connection. The proof follows similarly as in [GLS, Theorem 1.2.4] and is ommited.
@@
The existence a coupling form is a non-trivial question, related to the topology on the fibration. In the next result we recall a remarkable identity which relates the curvature of $\Gamma$ with the horizontal part of any coupling $\boldsymbol{\sigma}$ (see [GLS, Eq. (1.12)]). Recall that the curvature of an Ehresmann connection $\Gamma$ on $\curly{X}\to\curly{J}$ is the basic two-form $F_{\Gamma}\in\Omega^{2}(\curly{X},V\curly{X})$ defined by
@@
Assume the hypothesis of Theorem 3.2. Then,
@@
When the fibre $\curly D$} is compact and simply connected, the previous result implies the existence of a natural choice of coupling form $\boldsymbol{\sigma}$, given by imposing a ‘Gysin type condition’ (see [GLS, Th. 1.4.1]). The idea is that, under this hypothesis, the vertical vector field $F_{\Gamma}(v_{1},v_{2})$ is Hamiltonian, and hence it determines up to a constant the vertical variation of the function $\boldsymbol{\sigma}(v_{1}^{\Gamma},v_{2}^{\Gamma})$. More explicitly, provided that there exists a coherent choice of fibrewise Hamiltonian function $\mu(v_{1},v_{2})$ for $F_{\Gamma}(v_{1},v_{2})$, for any pair of horizontal vector fields $v_{1},v_{2}$, that is,
@@
(3.3) $\boldsymbol{\sigma}=\widehat{\boldsymbol{\omega}}(\Gamma,\Gamma)-\mu(v_{1},v_{2}).$
This is the case, for instance, if the connection $\Gamma$ has holonomy contained in a subgroup $K$ of Hamiltonian isommetries of the fibre $\curly D$}, and there is a $K$-equivariant moment map $\mu\colon{\curly D}\to\mathfrak{k}^{*}$.
@@
for a small coupling constant $0<\alpha\ll 1$.
In the present paper, we are interested in the study of Kähler fibrations satisfying the hypothesis of Theorem 3.2. We will find situations in which the pre-symplectic manifold $(\curly{X},\boldsymbol{\sigma})$ admits a Hamiltonian action of a real Lie group $\widetilde{\curly{G}}$, preserving the holomorphic fibration structure. This will have an impact in the structure of the group of symmetries, which will typically appear as a non-trivial extension
@@
of a (real) subgroup $\curly{H}$ of the holomorphic automorphisms of the base $\curly{J}$ by a subgroup of Kähler isometries $\curly{G}$ of the typical fibre $\curly D$}. Previous work on this type of reductions, in the context of gauge theory, can be found in [AGG, AGG3, AGGP, AGGPY].
3.2. Kähler connection and norm squared of the Higgs field

### v2

@@ 2.3. Harmonicity equations and hyperkähler structure
where $A$ is a connection on $E_{K}$ and $\psi\in\Omega^{1}(X,E_{K}(\mathfrak{k}))$. We thus have a one-to-one correspondence
(2.6) $\begin{array}[]{rcl}{\curly A}\times\Omega^{1}(\Sigma,E_{K}(\mathfrak{k}))&\longrightarrow&{\curly D}\\ (A,\psi)&\longmapsto&D:=A+i\psi,\end{array}$
@@
(2.7) $\begin{split}\mathbf{I}(a,\dot{\psi})&=(Ja,-J\dot{\psi}),\\ \mathbf{J}(a,\dot{\psi})&=(-\dot{\psi},a),\\ \mathbf{K}(a,\dot{\psi})&=(-J\dot{\psi},-Ja),\end{split}$
@@
(2.11) $\begin{split}F_{A}-\frac{1}{2}[\psi,\psi]&=0,\\ d_{A}\psi&=0,\\ d_{A}^{\ast}\psi&=0.\end{split}$
@@
The harmonicity condition for the reduction $h$ admits an interpretation in terms of the the holonomy representation $\rho:\pi_{1}(\Sigma)\to G$ corresponding to $D$. The reduction $h$ is equivalent to a $\pi_{1}(\Sigma)$-equivariant smooth map
$\tilde{h}:\tilde{\Sigma}\longrightarrow G/K,$
where $\tilde{\Sigma}$ is the universal covering of $\Sigma$. Here $\pi_{1}(\Sigma)$ acts on $\tilde{\Sigma}$ by Deck transformations, and via the representation $\rho$ on $G/K$. The symmetric space $G/K$ is equipped with a canonical $G$-invariant Riemannian metric determined by the Killing form $B$, and (2.11) is equivalent to the harmonicity of $\tilde{h}$ in the usual sense.
@@
The following theorem, due to Donaldson [Do1] for $G=\mathrm{SL}(2,\mathbb{C})$, and Corlette [Co] in general, is one of the main structural results of the theory.
@@
2.4. $\boldsymbol{G}$-Higgs bundles and Hitchin’s equations
@@
There are appropriate notions of stability, semistability and polystability and we can consider $\mathcal{M}^{Higgs}(G)$ to be the moduli space of isomorphism classes polystable $G$-Higgs bundles over $X$.
Let $h\in\Omega^{0}E(X,G/K)$ be a smooth reduction of structure group of $E$ to $K$ and let $E_{h}$ be the corresponding smooth principal $K$-bundle. Let $F_{h}\in\Omega^{2}(X,E_{h}(\mathfrak{k}))$ be the curvature of the unique connection compatible with $h$ and the holomorphic structure on $E$, defined by the Chern–Singer correspondence [Ch, Sin]. Here $E_{h}(\mathfrak{k})$ is the adjoint bundle of $E_{h}$.
@@
$\tau\colon\Omega^{1,0}(X,E(\mathfrak{g}))\longrightarrow\Omega^{0,1}(X,E(\mathfrak{g}))$
@@
A reduction $h$ of structure group of $E$ to $K$ satisfies Hitchin’s equation
@@
(2.12) $\begin{split}F_{A}-[\varphi,\tau(\varphi)]&=0,\\ \bar{\partial}_{A}\varphi&=0.\end{split}$
Here $\bar{\partial}_{A}$ is the $(0,1)$ part of the covariant derivative $d_{A}$ defined by $A$. This is of course the Dolbeault operator defined by the holomorphic structure $J_{A}$ on $E_{G}$ corresponding to $A$ via de Chern–Singer correspondence. The bundle $E_{G}$ is the smooth principal $G$-bundle obtained from $E_{K}$ by extension of structure group. The gauge group $\curly K$} of automorphisms of $E_{K}$ acts on the space of solutions, and the moduli space of solutions is
@@
$\mathcal{M}^{Higgs}(G)\cong\mathcal{M}^{\operatorname{Hit}}(G).$
To explain this correspondence we interpret the moduli space of $G$-Higgs bundles in terms of pairs $(J_{E},\varphi)$ consisting of a holomorphic structure $J_{E}$ on the smooth $G$-bundle $E_{G}$ obtained from $E_{K}$ by the extension of structure group, and $\varphi\in\Omega^{1,0}(X,E_{G}(\mathfrak{g}))$ satisfying $\bar{\partial}_{E}\varphi=0$. Such pairs are in correspondence with $G$-Higgs bundles $(E,\varphi)$, where $E$ is the holomorphic $G$-bundle defined by $J_{E}$ on $E_{G}$, and $\bar{\partial}_{E}\varphi=0$, that is, $\varphi\in H^{0}(X,E(\mathfrak{g})\otimes K_{X})$. The moduli space of polystable $G$-Higgs bundles $\mathcal{M}^{Higgs}(G)$ can now be identified with the orbit space
@@
2.5. Hitchin’s equations and hyperkähler structure
@@
As mentioned above, the Chern–Singer correspondence establishes an isomorphism
(2.13) $\begin{array}[]{rcl}{\curly A}&\longrightarrow&{\curly C}\\ A&\longmapsto&J_{A}.\end{array}$
@@
(2.14) $\begin{array}[]{rcl}\Omega^{0,1}(X,E_{G}(\mathfrak{g}))&\longrightarrow&\Omega^{1}(X,E_{K}(\mathfrak{k}))\\ \alpha&\longmapsto&a:=\alpha-\tau(\alpha).\end{array}$
@@
(2.15) $\begin{array}[]{rcl}{\curly A}&\longrightarrow&\Omega^{2}(X,E_{K}(\mathfrak{k}))\\ A&\longmapsto&F_{A}.\end{array}$
@@
(2.16) $\begin{array}[]{rcl}{\curly A}\times\Omega&\longrightarrow&\Omega^{2}(X,E_{K}(\mathfrak{k}))\\ (A,\varphi)&\longmapsto&F_{A}-[\varphi,\tau(\varphi)].\end{array}$
@@
$J_{1}(\alpha,\eta)$ $=(i\alpha,i\eta),$
$J_{2}(\alpha,\eta)$ $=(-i\tau(\eta),i\tau(\alpha)),$
$J_{3}(\alpha,\eta)$ $=(\tau(\eta),-\tau(\alpha)).$
These complex structures correspond to the complex structures $\mathbf{I},\mathbf{J}$ and $\mathbf{K}$ in (2.7) via the Chern–Singer correspondence.
@@
We thus have that ${{\boldsymbol{\mu}}}^{-1}({0})/{\curly K}$ is the moduli space $\mathcal{M}^{\operatorname{Hit}}(G)$ of solutions to Hitchin’s equations (2.12). In particular, if we consider the set ${{\boldsymbol{\mu}}}_{*}^{-1}({0})$ of irreducible solutions (equivalently, smooth) one has that
@@
is a hyperkähler manifold which, by Theorem 2.7, is homeomorphic to the subvariety of smooth points of the moduli space $\mathcal{M}^{Higgs}(G)$, consisting of stable and simple $G$-Higgs bundles on $E_{G}$.
@@ 2.6. Non-abelian Hodge correspondence
$\begin{array}[]{rcl}{\curly A}\times\Omega&\longrightarrow&{\curly D},\\ (A,\varphi)&\longmapsto&A-i\varphi+i\tau(\varphi),\end{array}$
@@
Now, Theorems 2.7 and 2.5 can be regarded as existence theorems, establishing the non-emptiness of the hyperkähler quotient, obtained by focusing on different complex structures. For Theorem 2.7 one gives a special status to the complex structure $J_{1}$. Combining the symplectic forms determined by $J_{2}$ and $J_{3}$ one has the $J_{1}$-holomorphic symplectic form $\omega_{c}=\omega_{2}+i\omega_{3}$ on ${\curly A}\times\Omega$. The gauge group $\curly{G}={\curly K}^{\mathbb{C}}$ acts on ${\curly A}\times\Omega$ preserving $\omega_{c}$. The symplectic quotient construction can also be extended to the holomorphic situation (see e.g. [Ko]) to obtain the holomorphic symplectic quotient $\{(J_{E},\varphi)\;:\;\bar{\partial}_{E}\varphi=0\}/\curly{G}$. What Theorem 2.7 says is that for a class $(J_{E},\varphi)]$ in this quotient to have a representative (unique up to $K$-gauge) satisfying $\mu_{1}=0$ it is necessary and sufficient that the pair $(J_{E},\varphi)$ be polystable. This identifies the hyperkähler quotient to the set of equivalence classes of polystable $G$-Higgs bundles on $E_{G}$. If one now takes $J_{2}$ on ${\curly A}\times\Omega$ or equivalently $\curly D$} with $\mathbf{J}$ and argues in a similar way, one gets Theorem 2.5 identifying the hyperkähler quotient to the set of equivalence classes of reductive flat connections on $E_{G}$.
@@ 3.1. Kähler fibrations
In this section we recall some basic aspects of the theory of Kähler fibrations, which we will use later. We follow closely [GLS, Mu].
A Kähler fibration is a holomorphic fibre bundle $\curly{X}\to\curly{J}$, with typical fibre a Kähler manifold $({\curly D},\omega_{\curly D})$, and such that transition functions between local holomorphic trivializations are contained in the group of Kähler isometries of $\curly D$}. Equivalently, $\curly{X}$ admits a smoothly varying Kähler structure on the fibres, that we shall denote $\widehat{\boldsymbol{\omega}}$. An Ehresmann connection $\Gamma$ on $\curly{X}$, given by a distribution of horizontal subspaces
@@
The following basic result shows that the existence of a closed $(1,1)$-form restricting on the fibres to $\widehat{\boldsymbol{\omega}}$ is indeed a sufficient condition for the existence of a Kähler connection. The proof follows similarly to [GLS, Theorem 1.2.4] and is omited.
@@
The existence of a coupling form is a non-trivial question, related to the topology of the fibration. In the next result we recall a remarkable identity which relates the curvature of $\Gamma$ with the horizontal part of any coupling $\boldsymbol{\sigma}$ (see [GLS, Eq. (1.12)]). Recall that the curvature of an Ehresmann connection $\Gamma$ on $\curly{X}\to\curly{J}$ is the basic 2-form $F_{\Gamma}\in\Omega^{2}(\curly{X},V\curly{X})$ defined by
@@
Assume the hypothesis of Theorem 3.2. Then
@@
When the fibre $\curly D$} is compact and simply connected, the previous result implies the existence of a natural choice of coupling form $\boldsymbol{\sigma}$, given by imposing a ‘Gysin type condition’ (see [GLS, Th. 1.4.1]). The idea is that, under this hypothesis, the vertical vector field $F_{\Gamma}(v_{1},v_{2})$ is Hamiltonian, and hence it determines up to a constant the vertical variation of the function $\boldsymbol{\sigma}(v_{1}^{\Gamma},v_{2}^{\Gamma})$. More explicitly, provided that there exists a coherent choice of a fibrewise Hamiltonian function $\mu(v_{1},v_{2})$ for $F_{\Gamma}(v_{1},v_{2})$, for any pair of horizontal vector fields $v_{1},v_{2}$, that is,
@@
(3.3) $\boldsymbol{\sigma}=\widehat{\boldsymbol{\omega}}(\Gamma,\Gamma)-\mu,$
where $\mu$ is regarded as a basic 2-form. This is the case, for instance, if the connection $\Gamma$ has holonomy contained in a subgroup $K$ of Hamiltonian isometries of the fibre $\curly D$}, and there is a $K$-equivariant moment map $\mu\colon{\curly D}\to\mathfrak{k}^{*}$.
@@
for a small ‘coupling constant’ $0<\alpha\ll 1$.
In the present paper, we are interested in the study of Kähler fibrations satisfying the hypothesis of Theorem 3.2. We will find situations in which the presymplectic manifold $(\curly{X},\boldsymbol{\sigma})$ admits a Hamiltonian action of a real Lie group $\widetilde{\curly{G}}$, preserving the holomorphic fibration structure. This will have an impact on the structure of the group of symmetries, which will typically appear as a non-trivial extension
@@
of a (real) subgroup $\curly{H}$ of the holomorphic automorphisms of the base $\curly{J}$ by a subgroup of Kähler isometries $\curly{G}$ of the typical fibre $\curly D$}. Previous work on this type of reductions, in the context of gauge theory, can be found in [AGG1, AGG3, AGGP, AGGPY].
3.2. The Kähler connection and the squared norm of the Higgs field